You are given three capacitors with the following capacitances: C1 = 4.42 mF, C2 = 9.57 mF, and C3 = 3.53 mF. What is the LARGEST equivalent capacitance you can build using these three capacitors?

ALSO, What is the SMALLEST equivalent capacitance you can assemble using these three capacitors?
Answer #1

when the capcitors are in parallel

Ceq=17.52 mF=sum

smallest when they are in series

1/Ceq=1/c1+1/c2+1/c3

Ceq=1.62 mC
Answer #2

Largest equivalent capacitance when connected in parallel =4.42+9.57+3.53=17.52 mF

Smallesr equivalent capacitance when connected in series =1.62 mF